### v1

@@ 4.3 LoRA Fine-tuning of Foundation Models
For the LoRA-based fine-tuning of the foundation models, each model was initialized with publicly available pre-trained weights and adapted using LoRA with rank 8, scaling factor 16, and dropout rate 0.3, applied to transformer attention and MLP layers. The classification head was re-initialized and trained jointly with LoRA modules, while the rest of the backbone remained frozen. This standardized protocol ensured fair and rigorous evaluation of LoRA-based adaptation across diverse foundation models on the AMi-Br benchmark.
@@
Model AMi-Br AtNorM-Br AtNorM-MD Balanced Acc. AUROC Balanced Acc. AUROC Balanced Acc. AUROC EfficientNetV2 $0.6488\pm 0.1265$ $0.7124\pm 0.1608$ $0.6457\pm 0.1055$ $0.7055\pm 0.1250$ $0.6034\pm 0.1166$ $0.6721\pm 0.1279$ Vision Transformer $0.7552\pm 0.0191$ $0.8634\pm 0.0104$ $0.7570\pm 0.0171$ $0.8478\pm 0.0189$ $0.7439\pm 0.0149$ $0.8396\pm 0.0213$ Swin Transformer $0.7491\pm 0.1061$ $0.8451\pm 0.0996$ $0.7213\pm 0.1116$ $0.8272\pm 0.0879$ $0.7226\pm 0.1048$ $0.8069\pm 0.1213$ UNI $0.6477\pm 0.0108$ $0.7171\pm 0.0080$ $0.6217\pm 0.0015$ $0.6774\pm 0.0136$ $0.5715\pm 0.0153$ $0.6241\pm 0.0222$ UNI2-h $0.6472\pm 0.0138$ $0.7087\pm 0.0107$ $0.6612\pm 0.0169$ $0.7296\pm 0.0115$ $0.5919\pm 0.0162$ $0.6297\pm 0.0173$ Virchow $0.6452\pm 0.0139$ $0.7010\pm 0.0077$ $0.6441\pm 0.0212$ $0.7037\pm 0.0107$ $0.5758\pm 0.0172$ $0.6175\pm 0.0227$ Virchow2 $0.6403\pm 0.0079$ $0.7040\pm 0.0052$ $0.6667\pm 0.0133$ $0.7201\pm 0.0060$ $0.5652\pm 0.0064$ $0.5878\pm 0.0211$ Prov-Gigapath $0.6091\pm 0.0140$ $0.6569\pm 0.0132$ $0.6005\pm 0.0091$ $0.6323\pm 0.0125$ $0.5601\pm 0.0153$ $0.5818\pm 0.0129$ H-Optimus-0 $0.6372\pm 0.0164$ $0.6926\pm 0.0062$ $0.6122\pm 0.0185$ $0.6903\pm 0.0217$ $0.5603\pm 0.0400$ $0.5935\pm 0.0423$ H-Optimus-1 $0.5918\pm 0.0165$ $0.6494\pm 0.0108$ $0.6139\pm 0.0195$ $0.6696\pm 0.0193$ $0.5915\pm 0.0308$ $0.6436\pm 0.0499$ UNI (LoRA) $0.7952\pm 0.0092$ $0.8839\pm 0.0059$ $0.7183\pm 0.0226$ $0.7979\pm 0.0142$ $0.7069\pm 0.0278$ $0.8222\pm 0.0224$ UNI2-h (LoRA) $0.7138\pm 0.0121$ $0.8153\pm 0.0211$ $0.7301\pm 0.0152$ $0.8228\pm 0.0143$ $0.6914\pm 0.0321$ $0.7616\pm 0.0415$ Virchow (LoRA) $0.7878\pm 0.0250$ $0.8891\pm 0.0150$ $\mathbf{0.7696}\pm 0.0198$ $0.8540\pm 0.0200$ $\mathbf{0.7705}\pm 0.0287$ $\mathbf{0.8641}\pm 0.0247$ Virchow2 (LoRA) $\mathbf{0.8135}\pm 0.0145$ $\mathbf{0.9026}\pm 0.0051$ $0.7632\pm 0.0190$ $\mathbf{0.8579}\pm 0.0117$ $0.7424\pm 0.0305$ $0.8503\pm 0.0171$ Prov-Gigapath (LoRA) $0.7602\pm 0.0113$ $0.8682\pm 0.0122$ $0.7263\pm 0.0296$ $0.8077\pm 0.0184$ $0.7007\pm 0.0228$ $0.8073\pm 0.0259$ H-Optimus-0 (LoRA) $0.5888\pm 0.0766$ $0.6159\pm 0.1276$ $0.5901\pm 0.0827$ $0.6224\pm 0.1220$ $0.5406\pm 0.0619$ $0.5826\pm 0.1263$ H-Optimus-1 (LoRA) $0.5908\pm 0.0909$ $0.6362\pm 0.1143$ $0.5699\pm 0.0752$ $0.6352\pm 0.1091$ $0.5617\pm 0.0874$ $0.6019\pm 0.1312$
@@ 5 Results
On the AMi-Br dataset, the best-performing model was Virchow2 with LoRA finetuning, achieving an average balanced accuracy of 0.8135 and an AUROC of 0.9026, outperforming all other models. Other LoRA-fine-tuned models such as UNI (0.7952 mean balanced accuracy) and Virchow (0.7878 balanced accuracy) also demonstrated strong performance, significantly improving over their respective linear probing counterparts. Among the end-to-end trained baseline models, the ViT achieved the highest balanced accuracy (0.7552) and AUROC (0.8634), outperforming both EfficientNetV2 and Swin Transformer. Foundation models with just linear probing lagged behind, with balanced accuracies around 0.59-0.64.
On the AtNorM-Br dataset, the trend was similar. Virchow with LoRA finetuning achieved the highest balanced accuracy of 0.7696 while Virchow2 with LoRA finetuning achieved the highest AUROC of 0.8579, followed by Virchow2 with LoRA finetuning (balanced accuracy: 0.7632). Other LoRA-fine-tuned models such as UNI2-h (0.7301) and UNI (0.7183) also showed consistent improvements over their linear probing variants. Among the baselines, the Vision Transformer maintained strong performance (balanced accuracy: 0.7570, AUROC: 0.8478), again outperforming EfficientNetV2 and Swin Transformer. Foundation models without LoRA adaptation performed moderately, with balanced accuracies between 0.60 and 0.66.
The AtNorM-MD dataset, which represents a distinct distribution shift, saw a general drop in performance across models. Virchow pretrained with LoRA performed the best, with a balanced accuracy of 0.7705 and AUROC of 0.8641. Other strong LoRA-fine-tuned models included Virchow2 (0.7424 balanced accuracy) and UNI (0.7069 balanced accuracy), all of which substantially outperformed their linear probing counterparts. The best baseline in this setting was again the Vision Transformer, achieving 0.7439 balanced accuracy and 0.8396 AUROC. In contrast, foundation models without LoRA adaptation performed less robustly. Across the board, LoRA fine-tuning consistently enhanced generalization performance under domain shift.
Across all the datasets, two of the models, H-Optimus-0 and H-Optimus-1, both fine-tuned with LoRA, consistently performed poorly, even less than their linear probing counterparts in balanced accuracy.
@@ 6 Discussion
Our evaluation confirmed that all foundation models exhibited some degree of generalization to the AMF identification task, but the extent of their effectiveness varied considerably. The application of LoRA fine-tuning revealed particularly striking performance differences across models.
The poor performance of H-Optimus-0 and H-Optimus-1 could be the result of a resolution mismatch, as the model expects input images trained at 0.5 microns/pixel resolution, whereas our dataset has images at 0.25 microns/pixel. In contrast, Virchow2 demonstrated superior performance, likely benefiting from training on higher resolution images using a 40x objective, making it more compatible with our dataset characteristics.
We also notice considerable drop in performance for the foundation models (both linear probing and LoRAs-fine-tuned) in the two external test datasets, as reflected by the balanced accuracy and AUROC scores. We attribute this to the domain shift between the datasets. For the AtNorM-Br dataset, the observed performance may partially also result from individually different criteria in the discrimination of AMFs, an effect that evens out when multiple raters are involved. As only a single annotator performed the labeling process for AtNorM-Br, this introduces potentially systematic label biases. Moreover, an image distribution shift cannot be ruled out, as the images were retrieved from different sources (labs, scanners, etc.), making the dataset more heterogeneous. For the AtNorM-MD dataset, where five annotators were involved, a systematic label bias is rather unlikely, however, the inclusion of multiple domains beyond breast cancer introduces a more pronounced image domain shift, encompassing a wider range of tissue types that pose significant challenges for model adaptation.
A significant limitation of our approach lies in training exclusively on the AMi-Br dataset with a relatively limited number of cases of only human breast cancer, restricting the models’ ability to learn the full spectrum of AMF representations across different tissue types, tumors, institutions, staining protocols, and scanners. Our results emphasize that automated AMF identification remains a highly challenging problem in computational pathology, requiring further methodological advances. Our new, multi-domain datasets, provided by this work, help further this cause.
@@
CAB, VW, and CK acknowledge the support from the Austrian Research Fund (FWF, project number: I 6555). SB, TC, RK, and MA acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, project number: 520330054). KB acknowledge support by the German Research Foundation (DFG) project 460333672 CRC1540 EBM.
@@
All data used in this study is public and can be found at this github repository: https://github.com/DeepMicroscopy/AMi-Br_Benchmark.

### v2

@@ 4.3 LoRA Fine-tuning of Foundation Models
For the LoRA-based fine-tuning of the foundation models, each model was initialized with publicly available pre-trained weights and adapted using LoRA with rank 8, scaling factor 16, and dropout rate 0.3, applied to both the transformer attention layers (i.e., query, key, value, and output projections) and the MLP components (i.e., feedforward layers fc1 and fc2) of the transformer blocks. The classification head was re-initialized and trained jointly with LoRA modules, while the rest of the backbone remained frozen. This standardized protocol ensured fair and rigorous evaluation of LoRA-based adaptation across diverse foundation models on the AMi-Br benchmark.
@@
Model AMi-Br AtNorM-Br AtNorM-MD Balanced Acc. AUROC Balanced Acc. AUROC Balanced Acc. AUROC EfficientNetV2 $0.7474\pm 0.0167$ $0.8315\pm 0.0168$ $0.7283\pm 0.0168$ $0.8110\pm 0.0084$ $0.6975\pm 0.0277$ $0.7604\pm 0.0217$ Vision Transformer $0.7899\pm 0.0268$ $0.8872\pm 0.0107$ $\textbf{0.7788}\pm 0.0229$ $\textbf{0.8710}\pm 0.0075$ $0.7494\pm 0.0342$ $0.8720\pm 0.0117$ Swin Transformer $0.8052\pm 0.0124$ $\textbf{0.9029}\pm 0.0112$ $0.7752\pm 0.0134$ $0.8683\pm 0.0113$ $\textbf{0.7723}\pm 0.0142$ $\textbf{0.8806}\pm 0.0087$ UNI $0.6339\pm 0.0189$ $0.6949\pm 0.0275$ $0.6406\pm 0.0339$ $0.7028\pm 0.0310$ $0.5510\pm 0.0123$ $0.5892\pm 0.0118$ UNI2-h $0.6220\pm 0.0095$ $0.6765\pm 0.0157$ $0.6482\pm 0.0214$ $0.7059\pm 0.0230$ $0.5773\pm 0.0212$ $0.6119\pm 0.0357$ Virchow $0.6029\pm 0.0132$ $0.6520\pm 0.0168$ $0.5969\pm 0.0155$ $0.6733\pm 0.0266$ $0.5623\pm 0.0235$ $0.6184\pm 0.0566$ Virchow2 $0.6089\pm 0.0192$ $0.6637\pm 0.0214$ $0.6236\pm 0.0296$ $0.6795\pm 0.0344$ $0.5694\pm 0.0173$ $0.6046\pm 0.0324$ Prov-Gigapath $0.6097\pm 0.0177$ $0.6435\pm 0.0188$ $0.5893\pm 0.0074$ $0.6187\pm 0.0193$ $0.5558\pm 0.0125$ $0.6039\pm 0.0244$ H-Optimus-0 $0.6519\pm 0.0089$ $0.7098\pm 0.0062$ $0.6191\pm 0.0156$ $0.7058\pm 0.0140$ $0.5541\pm 0.0323$ $0.5786\pm 0.0458$ H-Optimus-1 $0.6088\pm 0.0120$ $0.6705\pm 0.0094$ $0.6152\pm 0.0305$ $0.6868\pm 0.0195$ $0.6051\pm 0.0142$ $0.6596\pm 0.0294$ UNI (LoRA) $0.7952\pm 0.0092$ $0.8839\pm 0.0059$ $0.7183\pm 0.0226$ $0.7979\pm 0.0142$ $0.7069\pm 0.0278$ $0.8222\pm 0.0224$ UNI2-h (LoRA) $0.7138\pm 0.0121$ $0.8153\pm 0.0211$ $0.7301\pm 0.0152$ $0.8228\pm 0.0143$ $0.6914\pm 0.0321$ $0.7616\pm 0.0415$ Virchow (LoRA) $0.7878\pm 0.0250$ $0.8891\pm 0.0150$ $0.7696\pm 0.0198$ $0.8540\pm 0.0200$ $0.7705\pm 0.0287$ $0.8641\pm 0.0247$ Virchow2 (LoRA) $\mathbf{0.8135}\pm 0.0145$ $0.9026\pm 0.0051$ $0.7632\pm 0.0190$ $0.8579\pm 0.0117$ $0.7424\pm 0.0305$ $0.8503\pm 0.0171$ Prov-Gigapath (LoRA) $0.7602\pm 0.0113$ $0.8682\pm 0.0122$ $0.7263\pm 0.0296$ $0.8077\pm 0.0184$ $0.7007\pm 0.0228$ $0.8073\pm 0.0259$ H-Optimus-0 (LoRA) $0.7846\pm 0.0169$ $0.8721\pm 0.0148$ $0.7044\pm 0.0383$ $0.7921\pm 0.0271$ $0.6549\pm 0.0281$ $0.7879\pm 0.0163$ H-Optimus-1 (LoRA) $0.7762\pm 0.0110$ $0.8611\pm 0.0083$ $0.7658\pm 0.0298$ $0.8524\pm 0.0100$ $0.7396\pm 0.0202$ $0.8389\pm 0.0104$
@@ 5 Results
We now present the comparative performance of all models across the AMi-Br, AtNorM-Br, and AtNorM-MD datasets using balanced accuracy and AUROC as primary metrics. The full quantitative results are summarized in Table 3, where we report the mean and standard deviation on the hold-out datasets across five train-validation folds of the training part of AMi-Br. To complement the tabulated results, Figure 2 illustrates the corresponding receiver operating characteristic (ROC) curves for each dataset, comparing baseline models, foundation models with linear probing, and those fine-tuned using LoRA.
On the AMi-Br dataset, the best-performing model in terms of balanced accuracy was Virchow2 with LoRA finetuning, achieving an average balanced accuracy of 0.8135. While the highest AUROC of 0.9029 was obtained by the Swin Transformer model, Virchow2 closely followed with an AUROC of 0.9026, making it the most balanced performer overall. Other LoRA-fine-tuned models such as UNI (0.7952 mean balanced accuracy), Virchow (0.7878 mean balanced accuracy) and H-Optimus-0 (0.7846 mean balanced accuracy) also demonstrated strong performance, significantly improving over their respective linear probing counterparts. Among the end-to-end trained baseline models, the Swin Transformer achieved the highest balanced accuracy (0.8052) and AUROC (0.9029), outperforming both EfficientNetV2 and ViT. Foundation models with just linear probing lagged behind, with balanced accuracies ranging between 0.60-0.65.
On the AtNorM-Br dataset, the end-to-end trained ViT was the best performing model overall with a mean balanced accuracy of 0.7788 and AUROC 0.8710, while the Swin Transformer followed closely with a mean balanced accuracy of 0.7752 and an AUROC of 0.8683. Among the LoRA-fine-tuned foundation models, Virchow performed the best in terms of a mean balanced accuracy of 0.7696, followed by H-Optimus-1 (0.7658 mean balanced accuracy) and Virchow2 (0.7632 mean balanced accuracy) Other LoRA-fine-tuned models such as UNI2-h (0.7301 balanced accuracy), Prov-Gigapath (0.7263 balanced accuracy), UNI (0.7183 balanced accuracy) and H-Optimus-0 (0.7044 balanced accuracy) also showed consistent improvements over their linear probing variants. Foundation models with linear probing performed poorly, with balanced accuracies ranging between 0.59 and 0.65.
The AtNorM-MD dataset, which represents a distinct distribution shift, saw a general drop in performance across models. The end-to-end trained Swin Transformer performed the best, with a mean balanced accuracy of 0.7723 and an AUROC of 0.8806. Virchow pretrained with LoRA followed closely, with a balanced accuracy of 0.7705. Other LoRA-fine-tuned models also performed well, including Virchow2 (0.7424 balanced accuracy), H-Optimus-1 (0.7396 balanced accuracy), UNI (0.7069 balanced accuracy), Prov-Gigapath (0.7007 balanced accuracy), UNI2-h (0.6914 balanced accuracy) and H-Optimus-0 (0.6549 balanced accuracy), all of which substantially outperformed their linear probing counterparts. Among the end-to-end trained baselines, ViT followed closely after Swin Transformer, achieving 0.7494 balanced accuracy and 0.8720 AUROC. EfficientNetV2 achieved the least balanced accuracy of 0.6975 among the baselines. As with the other datasets, the foundation models without LoRA adaptation performed less robustly, with balanced accuracies ranging between 0.55 and 0.60.
@@ 6 Discussion
Our evaluation confirmed that all foundation models exhibited some degree of generalization to the AMF identification task, but the extent of their effectiveness varied considerably. The application of LoRA fine-tuning revealed particularly striking performance differences across models. Also, it is worthwhile to note that across nearly all datasets and performance metrics (balanced accuracy and AUROC), there consistently exists at least one end-to-end trained baseline model that outperforms both linear probing and LoRA-adapted foundation models.
We also notice a drop in performance for several settings, including both end-to-end trained models and foundation models (both linear probing and LoRAs-fine-tuned) in the two external test datasets, as reflected by the balanced accuracy and AUROC scores. We attribute this to the domain shift between the datasets. For the AtNorM-Br dataset, the observed performance may partially also be influenced by the reliance on a single expert involved in the classification of AMFs, potentially introducing systematic label biases and limiting generalizability, an effect that evens out when multiple raters are involved. Moreover, an image distribution shift cannot be ruled out, as the images were retrieved from different sources (labs, scanners, etc.), making the dataset more heterogeneous. For the AtNorM-MD dataset, where five annotators were involved, a systematic label bias is rather unlikely, however, the inclusion of multiple domains beyond breast cancer introduces a more pronounced image domain shift, encompassing a wider range of tissue types that pose significant challenges for model adaptation.
A significant limitation of our approach lies in training exclusively on the AMi-Br dataset with a limited number of cases of only human breast cancer, restricting the models’ ability to learn the full spectrum of AMF representations across different tissue types, tumors, institutions, staining protocols, and scanners. Our results emphasize that automated AMF identification remains a highly challenging problem in computational pathology, requiring further methodological advances. Our new, multi-domain datasets, provided by this work, help further this cause. Future work will address the integration of these classifier into a two-stage mitotic figure detection approach, and investigate the role of false positive / negative detections.
@@
CAB, VW, and CK acknowledge the support from the Austrian Research Fund (FWF, project number: I 6555). SB, TC, RK, and MA acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, project number: 520330054). KB acknowledges support by the German Research Foundation (DFG) project 460333672 CRC1540 EBM.
@@
All data used in this study is public and can be found at https://github.com/DeepMicroscopy/AMi-Br_Benchmark.